# Descriptive Image Quality Assessment in the Wild

### v2

@@ B.2 Details of Dataset Construction
As described in Sec. 3.3, for brief tasks, the questions and answers are all templated and randomly sampled from a pool. The questions of detailed reasoning tasks are also sampled from a pool. The constructed question pools and answer pools (if possible) of distortion identification, instant rating, assessment reasoning, and comparison reasoning tasks are illustrated in Tab. A6, Tab. A8, Tab. A6, and Tab. A8, respectively.
B.3 Dataset Statistics
Statistics of the response length in our DQ-495K dataset are detailed in Tab. A2. We provide statistics on both word count and string length. For the instant rating task, there is no distinction between single-distortion and multi-distortion cases. We also depict the word length distribution of detailed reasoning responses in Fig. A4.
Wordcloud map of our introduced DQ-495K dataset is given in Fig. A4. While plotting Fig. A4, we manually exclude “Image A” and “Image B”, since they are constant proper nouns across all texts. The most frequent words in our DQ-495K dataset (e.g., “overly high”, “color quantization”, “high contrast”, “high saturation”, and “detail”) are all highly relevant to the low-level properties and the visual quality of images.
Table A2: Response length statistics in our DQ-495K dataset, reported as word count / string length. For instant rating task, there is no distinction between single-distortion and multi-distortion cases.
Distortion Identification Assessment Reasoning Instant Rating Comparison Reasoning
Single-distortion 10.36 / 69.81 64.37 / 430.23 9.30 / 52.02 93.20 / 604.97
Multi-distortion 12.84 / 88.67 87.31 / 588.44 114.04 / 740.68
Figure A3: Word length distribution of detailed responses in our DQ-495K dataset.
Figure A4: Wordcloud map of our introduced DQ-495K dataset.
Appendix C Details of Model Setups
Model Architecture. DepictQA-Wild primarily adopts the architecture from DepictQA [71], structured as follows. Specifically, the input images and the question texts are first tokenized, then fused, finally processed by the Large Language Model (LLM) for response generation. (1) Tokenizing input images and question texts. We use a frozen CLIP pre-trained ViT-L/14 [45] as the image encoder to convert the input images into visual tokens. The question texts are tokenized into textual tokens using the SentencePiece tokenizer [21]. To bridge the different embedding spaces of visual and textual tokens, we implement a trainable image projector to map visual tokens into the textual space following [26, 84]. (2) Token fusion. We integrate the visual tokens into pre-defined positions within the textual tokens as token fusion. (3) Response generation using LLM. The fused tokens are fed into LLM to generate the final response. Here we mainly conduct experiments with Vicuna-v1.5-7B. Despite their capabilities, pre-trained LLMs typically do not perform well on IQA tasks without adjustments. Therefore, we employ LoRA [16], a fine-tuning technique that efficiently modifies a small subset of parameters within the LLM. Specifically, we apply LoRA to adjust the projection layers in all self-attention modules, following [16, 70]. This approach allows for targeted refinement of the model’s performance on IQA tasks without the need for extensive retraining.
Model setup. Since the CLIP pre-trained ViT-L/14 [45] encodes each $14\times 14$ patch to a visual token, the resolution of the input image should be integer multiples of 14. Therefore, we first pad the size of input images to integer multiples of 14 with zero-padding. Although we aim to retain the image resolution, constrained by computation resources, if the image is too large ($>$ 672), we bi-linearly resize the image with its original ratio retained. We encode the image patches into visual tokens using the CLIP pre-trained ViT-L/14 [45], with each token having a channel of 1024. The image projection layer projects these visual tokens to the hidden dimension of the LLM, which is 4096. In LoRA, the rank and scale factor are both set as 16. There are 32 attention layers in the LLM in total. In each attention layer, the projection weights of “query”, “key”, “value”, and “output” are adjusted using two delta parameters with the shape of $4096\times 16$ and $16\times 4096$, respectively.
Training and inference setup. In our experiments, we set the LoRA rank to 16. DepictQA-Wild is trained for 1 epoch with batch size 64. Adam optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, weight decay 0.001, and learning rate 0.0002 is used for training. The trainable parameters are 16M, constituting only 0.25% of the total parameters (7B). The model is trained on 8 GPUs (NVIDIA RTX A6000 48G). The training is completed in around 22 hours. During inference, the temperature is set to 0, since lots of predicted information (e.g., distortion, comparison result) need to be certain.
Appendix D More Results
[…]
The results are presented in Tab. A3. First, overall, retaining resolution is crucial for identifying images with better aspect ratio or higher resolution. Second, with down-sampling becomes severer (i.e., aspect ratio is from 0.95 to 0.5), the accuracy is improved since the quality drop is more significant. Third, for severe down-sampling (e.g., aspect ratio is 0.5) where the quality degradation is quite obviously, retaining resolution or just resizing both images to a larger size both perform well ($\leq$ 99.0). Finally, however, for relatively slight down-sampling (e.g., aspect ratio is from 0.75 to 0.95), the performance of retaining resolution is stably superior than resizing.
Table A3: Retaining resolution during both training and inference is important to identify images with better aspect ratio or higher resolution.
[…]
Training Inference
73.0 99.0 93.5 91.7 83.8 77.2 71.2
85.6 99.8 99.4 99.0 95.9 94.8 89.4
98.8 99.9 99.6 99.3 99.1 96.8 97.0
Figure A5: Our estimated confidence scores are high correlated to the model performance on (a) distortion identification and (b) (c) instant rating tasks on different benchmarks in both full-reference and non-reference settings.
Ablation study on confidence estimation. We further examine the correlation between model performance and estimated confidence scores on a wider range of benchmarks. The results are illustrated in Fig. A5. The performance of our model is consistently enhanced as the confidence interval increases, validating the effectiveness of our confidence estimation.
Figure A6: An example of the model-restored image.
Details of quality comparison on model-processed images. We consider five common image restoration tasks: super-resolution, denoising, JPEG compression artifact removal, motion deblurring, and defocus deblurring. For each task, we collect three to four cutting-edged models in recent years (listed in Tab. A4), apply them to a correspondingly degraded image, and then manually rank the resultant model-processed images. To find the image considered best by VLMs, we linearly scan the candidates and compare them in pairs. As VLMs’ results are not deterministic and may be sensitive to the presentation order of images, we repeat the linear scan 10 times and randomly shuffle the scan order each time.
We have shown that our DepictQA-Wild can consistently find the near-optimal candidate compared to GPT-4V and scored-based methods. A highlight is that our DepictQA-Wild generalizes well on these out-of-distribution (OOD) model-processed images. For example, the image in Fig. A6 is restored from a noisy image. There is still remnant noise, which is somewhat strange. For such an OOD image, our DepictQA-Wild correctly recognizes it to be inferior, but MANIQA [66], MUSIQ [20], and NIQE [37] consider it as the best of the four candidates.
[t]
Table A4: Image restoration models used in quality comparison on model-processed images.
Image restoration task Image restoration models
Super-resolution SwinIR [23], HAT [7], X-Restormer [6]
Denoising SwinIR [23]1, MPRNet [73], Restormer [74], X-Restormer [6]
JPEG compression artifact removal SwinIR [23]2, FBCNN ($q$=90) [18]3, FBCNN (blind) [18]4
Motion deblurring MAXIM [57], MPRNet [73], Restormer [74]
Defocus deblurring DRBNet [47], IFAN [22], Restormer [74]
  • 1
    The variant trained on the noise level 15.
  • 2
    The variant trained on the quality factor 40.
  • 3
    The variant trained on the quality factor 90.
  • 4
    The variant blind to the quality factor.
@@ Appendix D More Results
More Qualitative Results. More qualitative results of assessment reasoning, comparison reasoning, and assessment on web-downloaded images are presented in Fig. A7, Fig. A8, Fig. A9, Fig. A10, and Fig. A11. DepictQA-Wild could accurately identify distortions, analyze their impacts on the display of image contents, then weigh the advantages and disadvantages of different aspects, and finally draw a final conclusion (e.g., overall quality, comparison results).
Figure A7: Qualitative results on assessment reasoning task in the full-reference setting. The two images from top to down are the reference image and evaluated image, respectively.
Figure A8: Qualitative results on assessment reasoning task in the non-reference setting.
Figure A9: Qualitative results on comparison reasoning task in the full-reference setting. The three images from left to right are the reference image, Image A, and Image B, respectively.
Figure A10: Qualitative results on comparison reasoning task in the non-reference setting. The two images from top to down are Image A and Image B, respectively.
Figure A11: Qualitative results on assessing web-downloaded images.
Table A5: Question pool and answer pool of distortion identification task.
@@ 24 Q: What distortion(s) most notably affect the clarity of the evaluated image?
Table A6: Question pool of assessment reasoning task.
@@ 18 How would you rate the overall quality of the image, and why?
Table A7: Question pool and answer pool of instant rating task.
@@ 19 Q: Which image, Image A or Image B, do you think displays better quality when compared?
Table A8: Question pool of comparison reasoning task.

### v3

@@ VI Conclusions and Limitations
As stated in Sec. III-C, for brief tasks, the questions and answers are templated and sampled from a pool. The questions of detailed tasks are also sampled from a pool. The question pools and answer pools (if possible) of distortion identification, instant rating, assessment reasoning, and comparison reasoning tasks are given in Tab. S5, Tab. S7, Tab. S5, and Tab. S7, respectively.
-B More Ablation Studies
Influence of comparison numbers. We calculate the win rate of one image over other compared images as the quality score. Here the compared images are selected by round robin for a small number, and random sampling for a large number. For the SPAQ dataset, the number of possible compared images is quite large, thus we adopt the random sampling strategy. The influence of comparison numbers is investigated in Tab. S3. It is shown that the comparison number could be reduced significantly without large performance degradation. In the most extreme cases (i.e., the comparison number is 1 or 2), we use the estimated confidence as weights to calculate the win rate as quality score. Otherwise, the values of the win rate are too discrete (i.e., the values of the win rate can only be 0 or 1 when the comparison number is 1). The results of our DepictQA-Wild are still reasonable in such extreme cases. Considering that the random sampling may bring large randomness or variance, we average the results with 5 random runs for small comparison numbers (i.e., << 10). Although the comparison number is small and the sampling process is random, our method is still very stable with relatively small standard deviations in Tab. S3.
TABLE S3: Influence of comparison numbers per image on SPAQ dataset with SRCC and PLCC metrics. “(KONIQ)” means the model is trained on KONIQ dataset, which is also an in-the-wild IQA dataset. For small comparison numbers (<< 10), we average the results with 5 random runs.
Comparison Numbers 100 50 25 10 5 2 1
DepictQA-Wild (Original) SRCC 0.835 0.832 0.826 0.806 0.731 $\pm$ 0.006 0.647 $\pm$ 0.009 0.577 $\pm$ 0.015
PLCC 0.841 0.837 0.832 0.810 0.735 $\pm$ 0.006 0.639 $\pm$ 0.009 0.537 $\pm$ 0.015
DepictQA-Wild (KONIQ) SRCC 0.859 0.854 0.850 0.830 0.756 $\pm$ 0.006 0.664 $\pm$ 0.011 0.598 $\pm$ 0.013
PLCC 0.861 0.858 0.852 0.833 0.757 $\pm$ 0.007 0.652 $\pm$ 0.011 0.546 $\pm$ 0.015
-C More Qualitative Results
@@
More qualitative results of assessment reasoning, comparison reasoning, and assessment on web-downloaded images are presented in Fig. S4, Fig. S5, Fig. S6, Fig. S7, and Fig. S8, respectively. Our DepictQA-Wild could accurately identify distortions, analyze their impacts on the display of image contents, then weigh the advantages and disadvantages of different aspects, and finally draw a final conclusion.
Figure S2: Humans usually identify at most two distortions (blue) when three (purple) are applied.
TABLE S4: Question pool and answer pool of distortion identification task.
@@ 24 Q: What distortion(s) most notably affect the clarity of the evaluated image?
TABLE S5: Question pool of assessment reasoning task.
@@ 18 How would you rate the overall quality of the image, and why?
TABLE S6: Question pool and answer pool of instant rating task.
@@ 19 Q: Which image, Image A or Image B, do you think displays better quality when compared?
TABLE S7: Question pool of comparison reasoning task.
@@ 11 Can you evaluate the overall quality in both Image A and Image B, and explain which one is superior?
Figure S3: Distortion examples of our distortion design. We showcase one example for each distortion implementation. The reference image is depicted in Fig. S1.
Figure S4: Qualitative results on assessment reasoning task in the full-reference setting. The two images from top to down are the reference image and evaluated image, respectively.
Figure S5: Qualitative results on assessment reasoning task in the non-reference setting.
Figure S6: Qualitative results on comparison reasoning task in the full-reference setting. The three images from left to right are the reference image, Image A, and Image B, respectively.
Figure S7: Qualitative results on comparison reasoning task in the non-reference setting. The two images from top to down are Image A and Image B, respectively.
Figure S8: Qualitative results on assessing web-downloaded images.